OpenStudy (anonymous):

Solve the equation in the interval [0,\, 2 \pi]. If there is more than one solution write them separated by commas. Hint: To solve this problem you will have to use the quadratic formula, inverse trigonometric functions and the symmetry of the unit circle. Also, you should look at the graph of the tangent function over the interval [0,\, 2 \pi]. (\tan x) ^2 - 21 \tan(x) - 22 = 0

OpenStudy (anonymous):

(\tan x) ^2 - 21 \tan(x) - 22 = 0 what is this?

OpenStudy (anonymous):

tan(x)^2-21tan(x)-22

OpenStudy (anonymous):

I know how to solve it but i cant get the points. I need the values of x in radian form

OpenStudy (turingtest):

let \(u=\tan x\) we then have\[\tan^2x-21\tan x-22=0\]\[u^2-21u-22=0\]solve for u, then sub back in and use inverse tangent to find x

OpenStudy (anonymous):

I know that U=22, -1. Which means tan (x)=22, tan (x)=-1. After this is where I need help with

OpenStudy (waleed_imtiaz):

If I use the quadratic formula to solve this without assuming u=tanx.... just use the formula as it is..... I Have got... Tanx=32.5 and tanx=-1...... now just use inverse tan... U will get the answer......